In a raffle, you can pay $10 to select a three-digit number from 000 through 199. If you select the same sequence of three digits that are drawn, you win and collect $100 otherwise you get nothing.

a.   How many different selections are possible?

b.   What is the probability of winning?

c.   If you win, what is your net profit?

d.   Complete the table below to find the expected value of this raffle

Event

X P(X) X*P(X)

Win(net gain)

Lose

E=
Answer #1

a) total different selections =200 (from 000 to 199 total numbers are 200)

b)probability of winning =1/200 =0.005

c)P(net profit) =100-10 =90

d)

x p(x) x*P(x)
win(net gain) 90 0.005 0.45
lose -10 0.995 -9.95
total -9.5